# Coding for Gaussian Two-Way Channels: Linear and Learning-Based Approaches

### v1

@@ III-A Signal Model for Linear Coding
III-B Optimal Linear Decoding
Since the decoding is conducted at each of the users independently, we can use a similar technique used in feedback-enabled GOWCs to find the optimal linear decoding scheme [10, 11]. After the $N$ channel uses, each user aims to estimate the message of the other user. We first consider the decoding process at User 1. User 1 estimates $m_{2}$ with the receive signal ${\bf y}_{1}$ by using a linear combining vector ${\bf w}_{2}\in\mathbb{R}^{N\times 1}$. We note that ${\bf y}_{1}={\bf x}_{2}+{\bf n}_{2}={\bf F}_{2}{\bf g}_{1}m_{1}+({\bf I}+{\bf F}_{2}{\bf F}_{1}){\bf g}_{2}m_{2}+{\bf F}_{2}({\bf I}+{\bf F}_{1}{\bf F}_{2}){\bf n}_{1}+({\bf I}+{\bf F}_{2}{\bf F}_{1}){\bf n}_{2}$, where ${\bf x}_{2}$ is given in (15). Rather than using ${\bf y}_{1}$ directly, we consider the following two pre-processing stages: (i) User 1 obtains ${{\bf z}_{1}}={\bf y}_{1}-{\bf F}_{2}{\bf g}_{1}m_{1}=({\bf I}+{\bf F}_{2}{\bf F}_{1}){\bf g}_{2}m_{2}+({\bf I}+{\bf F}_{2}{\bf F}_{1}){\bf F}_{2}{\bf n}_{1}+({\bf I}+{\bf F}_{2}{\bf F}_{1}){\bf n}_{2}$ by subtracting its message contribution, ${\bf F}_{2}{\bf g}_{1}m_{1}$, from ${\bf y}_{1}$; (ii) User 1 subsequently defines ${\tilde{\bf y}_{1}}=({\bf I}+{\bf F}_{2}{\bf F}_{1})^{-1}{\bf z}_{1}={\bf g}_{2}m_{2}+{\bf F}_{2}{\bf n}_{1}+{\bf n}_{2}$. Using ${{\bf y}_{1}}$ is equivalent to using ${\tilde{\bf y}_{1}}$ in terms of estimating $m_{2}$, since ${\tilde{\bf y}_{1}}$ is a linear function of ${{\bf y}_{1}}$. Using the result of pre-processing, User 1 obtains the message estimate as ${\hat{m}}_{2}={\bf w}_{2}^{\top}\tilde{\bf y}_{1}$. Since $\tilde{\bf y}_{1}={\bf g}_{2}m_{2}+{\bf F}_{2}{\bf n}_{1}+{\bf n}_{2}$, we represent ${\hat{m}}_{2}={\bf w}_{2}^{\top}{\bf g}_{2}m_{2}+{\bf w}_{2}^{\top}({\bf F}_{2}{\bf n}_{1}+{\bf n}_{2})$, where the first term is the message portion, while the second terms denotes the effective noise. The effective noise, ${\bf w}_{2}^{\top}({\bf F}_{2}{\bf n}_{1}+{\bf n}_{2})$, still follows a Gaussian distribution due to the linear processing of the Gaussian noise variables, ${\bf n}_{1}$ and ${\bf n}_{2}$. The distribution of the effective noise is given by $\mathcal{N}(0,{{\bf w}_{2}^{\top}{\bf Q}_{2}{\bf w}_{2}})$, where
${\bf Q}_{2}$ $={\bf F}_{2}{\bf F}_{2}^{\top}\sigma_{1}^{2}+\sigma_{2}^{2}{\bf I}.$ (18)
The SNR used to estimate $m_{2}$ is correspondingly defined by
@@ III-B Optimal Linear Decoding
Similarly, we consider that User 2 estimates $m_{1}$ using the receive signal ${\bf y}_{2}$ via a linear combining vector ${\bf w}_{1}\in\mathbb{R}^{N\times 1}$. In pre-processing, User 2 subtracts its message contribution, ${\bf F}_{1}{\bf g}_{2}m_{2}$, from ${\bf y}_{2}$ in (13), and obtains $\tilde{\bf y}_{2}={\bf y}_{2}-{\bf F}_{1}{\bf g}_{2}m_{2}={\bf g}_{1}m_{1}+({\bf I}+{\bf F}_{1}{\bf F}_{2}){\bf n}_{1}+{\bf F}_{1}{\bf n}_{2}$. User 2 then obtains the message estimate as ${\hat{m}}_{1}={\bf w}_{1}^{\top}\tilde{\bf y}_{2}$. By using the expression of $\tilde{\bf y}_{2}$ obtained above, we obtain ${\hat{m}}_{1}={\bf w}_{1}^{\top}{\bf g}_{1}m_{1}+{\bf w}_{1}^{\top}\big{(}({\bf I}+{\bf F}_{1}{\bf F}_{2}){\bf n}_{1}+{\bf F}_{1}{\bf n}_{2}\big{)}$, where the first and second terms are the message portion and the effective noise, respectively. The effective noise term, ${\bf w}_{1}^{\top}\big{(}({\bf I}+{\bf F}_{1}{\bf F}_{2}){\bf n}_{1}+{\bf F}_{1}{\bf n}_{2}\big{)}$, follows the Gaussian distribution with $\mathcal{N}(0,{{\bf w}_{1}^{\top}{\bf Q}_{1}{\bf w}_{1}})$, where
${\bf Q}_{1}$ $=({\bf I}+{\bf F}_{1}{\bf F}_{2})({\bf I}+{\bf F}_{1}{\bf F}_{2})^{\top}\sigma_{1}^{2}+{\bf F}_{1}{\bf F}_{1}^{\top}\sigma_{2}^{2}.$ (20)
The corresponding SNR to estimate $m_{1}$ is defined by
@@
Given ${\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, and ${\bf F}_{2}$, the optimal combining vector (that minimizes the error probability for message estimation) is obtained by maximizing the SNRs in (19) and (21[19, 10, 11], given by
${\bf w}^{\star}_{i}=\frac{{\bf Q}_{i}^{-1}{\bf g}_{i}}{{\bf g}_{i}^{\top}{{\bf Q}_{i}}^{-1}{\bf g}_{i}},\quad i\in\{1,2\}.$ (22)
Plugging (22) in (19) and (21), we obtain the following SNR expression
@@
The optimal combining vector for decoding in (22) is represented as a function of the encoding schemes of the users, i.e., ${\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, and ${\bf F}_{2}$. Accordingly, the SNR expression in (23) is also a function of the encoding schemes. Thus, the joint design of the encoding and decoding schemes is equivalent to the design of the encoding schemes only with the derived form of SNRs in (23).
@@ III-C Conversion of Optimization Problem
subject to $\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]}\leq NP,\quad\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\leq NP,$ (25)
where $\text{SNR}_{i}$ and ${\bf x}_{i}$, $i\in\{1,2\}$, are functions of the variables, ${\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, and ${\bf F}_{2}$, and the expectation is taken over the distributions of ${m}_{1}$, ${m}_{2}$, ${\bf n}_{1}$, and ${\bf n}_{2}$. However, it is intractable to directly solve the optimization problem in (25) due to (i) the coupling between the encoding schemes and the SNR expression in (23) and (ii) the power constraints in (25). We thus propose to transform it into a more tractable optimization problem through several steps as explained below. We start by stating the following remark.
@@
[…]
We further discuss the validity of Remark 1 within our linear coding framework in Appendix B. Based on the relationship between sum-error and power in Remark 1, we introduce an alternative optimization problem to (25), which aims to minimize the maximum transmit power, given by
$\underset{{\bf g}_{1},{\bf F}_{1},{\bf g}_{2},{\bf F}_{2}}{\text{minimize}}$ $\bar{P}$
@@
$\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]}\leq N\bar{P},\quad\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\leq N\bar{P},$ (26)
where $\bar{P}$ denotes the objective value which is distinguished from the average power constraint value $P$. We consider that $w$ is given, i.e., not a variable, in solving (26), and thus the two problems in (25) and (26) are not equivalent. However, the optimal solutions for (25) can be obtained by solving (26) through adjusting the value of $w$ based on Lemma 1; if the obtained objective value $\bar{P}$ in (26) is smaller than $P$, we keep decreasing the value of $\omega$ and solve the above problem until $\bar{P}$ reaches $P$. However, it is still challenging to solve (26) under the sum-error equality constraint, since there may exist multiple pairs of $\text{SNR}_{1}$ and $\text{SNR}_{2}$ that satisfy the constraint. Therefore, we consider individual constraints on $\text{SNR}_{1}$ and $\text{SNR}_{2}$ for optimization rather than the sum-error constraint. We subsequently reformulate (26) and obtain
$\underset{{\bf g}_{1},{\bf F}_{1},{\bf g}_{2},{\bf F}_{2}}{\text{minimize}}$ $\bar{P}$
subject to $\text{SNR}_{1}=\eta_{1},\quad\text{SNR}_{2}=\eta_{2}.$
$\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]}\leq N\bar{P},\quad\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\leq N\bar{P},$ (27)
where $\eta_{i}\in\mathbb{R}^{+}$ is the target SNR for the message $m_{i}$, $i\in\{1,2\}$. The two problems (26) and (27) become equivalent when $\mathcal{E}_{1}(\eta_{1})+\mathcal{E}_{2}(\eta_{2})=\omega$. To achieve equivalence in obtaining optimal solutions between the original problem in (25) and the modified problem in (27), we solve (27) for all feasible values of $\eta_{1}$ and $\eta_{2}$ and find the best solutions that minimize $\mathcal{E}_{1}(\eta_{1})+\mathcal{E}_{2}(\eta_{2})$ while satisfying $\bar{P}\leq P$.
If $\bar{P}$ is an optimal value of the problem in (27), at least one of the power constraints in (27) should be satisfied in equality, i.e., $\bar{P}=\max\{\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]},\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\}/N$. If $\bar{P}\neq\max\{\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]},$ $\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\}/N$, $\bar{P}$ is not an optimal solution since $\max\{\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]},\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\}/N$ can be the smaller objective value in (27). This insight leads us to reformulate (27) equivalently to
$\underset{{\bf g}_{1},{\bf F}_{1},{\bf g}_{2},{\bf F}_{2}}{\text{minimize}}$ ${\max}\big{\{}\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]},\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\big{\}}$
@@
We derive the following lemma, which is useful in solving (28).

### v2

@@ III-A Signal Model for Linear Coding
III-B Maximum Likelihood Decoding
Although encoding at the users is interrelated, decoding is an independent process at each user given the encoding schemes. Therefore, given the users’ encoding schemes, i.e., ${\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, and ${\bf F}_{2}$, the optimal decoding strategy to minimize the sum-error is for each user to conduct its own optimal decoding. The optimal decoding for each user is to select the message that maximizes the a posteriori probability among all possible messages. Since we assume that all messages are equi-probable, the optimal decoding is the maximum likelihood decoding [9]. In this subsection, we derive the maximum likelihood decoding for each user to estimate the message of the other user.
We first consider the decoding process at User 1. User 1 estimates $m_{2}$ given the receive signal ${\bf y}_{1}$ and its own message $m_{1}$ based on maximum likelihood, i.e., $\hat{m}_{2}=\underset{m_{2}\in\mathcal{M}_{2}}{\arg\max}\;\text{Pr}({\bf y}_{1}|m_{1},m_{2})$. By using the expression of ${\bf x}_{2}$ in (15), ${\bf y}_{1}$ is expressed as
${\bf y}_{1}={\bf x}_{2}+{\bf n}_{2}={\bf F}_{2}{\bf g}_{1}m_{1}+({\bf I}+{\bf F}_{2}{\bf F}_{1}){\bf g}_{2}m_{2}+({\bf I}+{\bf F}_{2}{\bf F}_{1}){\bf F}_{2}{\bf n}_{1}+({\bf I}+{\bf F}_{2}{\bf F}_{1}){\bf n}_{2}.$ (18)
We first define ${\bf r}_{1}$ with a linear transformation of ${\bf y}_{1}$ as
${\bf r}_{1}$ $=({\bf I}+{\bf F}_{2}{\bf F}_{1})^{-1}({\bf y}_{1}-{\bf F}_{2}{\bf g}_{1}m_{1})$
$={\bf g}_{2}m_{2}+{\bf F}_{2}{\bf n}_{1}+{\bf n}_{2},$ (19)
where $({\bf I}+{\bf F}_{2}{\bf F}_{1})$ is always invertible for any strictly lower triangular matrices ${\bf F}_{1}$ and ${\bf F}_{2}$, since $({\bf I}+{\bf F}_{2}{\bf F}_{1})$ has all 1’s on its diagonal entries. We note that ${\bf r}_{1}$ is a sufficient statistic for estimating $m_{2}$ given $m_{1}$ since, by definition, the conditional distribution of ${\bf y}_{1}$ given ${\bf r}_{1}$ does not depend on $m_{2}$. In other words, ${\bf r}_{1}$ has all the information about $m_{2}$. Then, we have $\hat{m}_{2}=\underset{m_{2}\in\mathcal{M}_{2}}{\arg\max}\;\;\text{Pr}({\bf y}_{1}|m_{1},m_{2})=\underset{m_{2}\in\mathcal{M}_{2}}{\arg\max}\;\;\text{Pr}({\bf r}_{1}|m_{1},m_{2})$. Note that ${\bf r}_{1}$ is conditionally Gaussian with conditional mean and conditional covariance matrix, given by
[…]
Thus, we obtain
$\hat{m}_{2}$ $=\underset{m_{2}\in\mathcal{M}_{2}}{\arg\max}\;\;\text{Pr}({\bf r}_{1}|m_{1},m_{2})$
$=\underset{m_{2}\in\mathcal{M}_{2}}{\arg\min}\;\;({\bf r}_{1}-{\bf g}_{2}m_{2})^{\top}{\bf Q}_{2}^{-1}({\bf r}_{1}-{\bf g}_{2}m_{2}).$ (22)
Note that ${\bf Q}_{2}$ defined in (21) is invertible, since ${\bf F}_{2}{\bf F}_{2}^{\top}$ is positive semi-definite (PSD), and thus ${\bf F}_{2}{\bf F}_{2}^{\top}\sigma_{1}^{2}+\sigma_{2}^{2}{\bf I}$ has all eigenvalues greater than or equal to $\sigma_{2}^{2}$, ensuring ${\bf Q}_{2}$ has full rank. The quadratic form in (22) can be equivalently rewritten as
$({\bf r}_{1}-{\bf g}_{2}\tilde{m}_{2})^{\top}{\bf Q}_{2}^{-1}({\bf r}_{1}-{\bf g}_{2}\tilde{m}_{2})+(m_{2}-\tilde{m}_{2})^{2}{\bf g}_{2}^{\top}{\bf Q}_{2}^{-1}{\bf g}_{2},$ (23)
by defining
$\tilde{m}_{2}=\frac{{\bf g}_{2}^{\top}{\bf Q}_{2}^{-1}}{{\bf g}_{2}^{\top}{\bf Q}_{2}^{-1}{\bf g}_{2}}{\bf r}_{1}:={\bf w}_{2}^{\top}{\bf r}_{1}\in\mathbb{R}.$ (24)
Note that $\tilde{m}_{2}$ in (24) is the minimum variance unbiased (MVU) estimate of $m_{2}$, and
${\bf w}_{2}=\frac{{\bf Q}_{2}^{-1}{\bf g}_{2}}{{\bf g}_{2}^{\top}{\bf Q}_{2}^{-1}{\bf g}_{2}}$ (25)
is the MVU estimator as a combining vector. Using (23), we rewrite (22) as
$\hat{m}_{2}=\underset{m_{2}\in\mathcal{M}_{2}}{\arg\min}\;\;(m_{2}-\tilde{m}_{2})^{2}.$ (26)
Thus, the optimal decision for User $1$ is to select $m_{2}\in\mathcal{M}_{2}$ that is closest to the MVU estimate $\tilde{m}_{2}\in\mathbb{R}$.
Next, we formulate the receive signal-to-noise (SNR) in estimating $m_{2}$ at User 1. By putting ${\bf r}_{1}$ (in (19)) into (24), we can rewrite (24) as
$\tilde{m}_{2}={\bf w}_{2}^{\top}{\bf g}_{2}m_{2}+{\bf w}_{2}^{\top}({\bf F}_{2}{\bf n}_{1}+{\bf n}_{2}).$ (27)
Since $\tilde{m}_{2}$ is also a sufficient statistic for estimating $m_{2}$, by using (27), we can define the receive SNR in estimating $m_{2}$ as
@@ III-B Maximum Likelihood Decoding
Next, we consider the decoding process at User 2. User 2 estimates $m_{1}$ given the receive signal ${\bf y}_{2}$ and its own message $m_{2}$ based on maximum likelihood. i.e., $\hat{m}_{1}=\underset{m_{1}\in\mathcal{M}_{1}}{\arg\max}\;\text{Pr}({\bf y}_{2}|m_{1},m_{2})$. Recall ${\bf y}_{2}={\bf g}_{1}m_{1}+{\bf F}_{1}{\bf g}_{2}m_{2}+({\bf I}+{\bf F}_{1}{\bf F}_{2}){\bf n}_{1}+{\bf F}_{1}{\bf n}_{2}$ in (13). We first define ${\bf r}_{2}$ with a linear transformation of ${\bf y}_{2}$ as
${\bf r}_{2}$ $={\bf y}_{2}-{\bf F}_{1}{\bf g}_{2}m_{2}.$
$={\bf g}_{1}m_{1}+({\bf I}+{\bf F}_{1}{\bf F}_{2}){\bf n}_{1}+{\bf F}_{1}{\bf n}_{2}.$ (29)
We note that ${\bf r}_{2}$ is a sufficient statistic for estimating $m_{1}$ given $m_{2}$, since, by definition, the conditional distribution of ${\bf y}_{2}$ given ${\bf r}_{2}$ does not depend on $m_{1}$. Then, we obtain $\hat{m}_{1}=\underset{m_{1}\in\mathcal{M}_{1}}{\arg\max}\;\;\text{Pr}({\bf y}_{2}|m_{1},m_{2})=\underset{m_{1}\in\mathcal{M}_{1}}{\arg\max}\;\;\text{Pr}({\bf r}_{2}|m_{1},m_{2})$. Note that ${\bf r}_{2}$ is conditionally Gaussian with conditional mean and conditional covariance matrix, given by
[…]
Thus, we obtain
$\hat{m}_{1}$ $=\underset{m_{1}\in\mathcal{M}_{1}}{\arg\max}\;\;\text{Pr}({\bf r}_{2}|m_{1},m_{2})$
$=\underset{m_{1}\in\mathcal{M}_{1}}{\arg\min}\;\;({\bf r}_{2}-{\bf g}_{1}m_{1})^{\top}{\bf Q}_{1}^{-1}({\bf r}_{2}-{\bf g}_{1}m_{1}).$ (32)
Note that ${\bf Q}_{1}$ defined in (31) is invertible, since (i) $({\bf I}+{\bf F}_{1}{\bf F}_{2})({\bf I}+{\bf F}_{1}{\bf F}_{2})^{\top}\sigma_{1}^{2}$ has all eigenvalues greater than or equal to $\sigma_{1}^{2}$ for strictly lower triangular matrices ${\bf F}_{1}$ and ${\bf F}_{2}$, and (ii) ${\bf F}_{1}{\bf F}_{1}^{\top}$ is PSD, and thus the summation of (i) and (ii) has all eigenvalues greater than or equal to $\sigma_{1}^{2}$, ensuring ${\bf Q}_{1}$ has full rank. The quadratic form in (29) is equivalently rewritten as
$({\bf r}_{2}-{\bf g}_{1}\tilde{m}_{1})^{\top}{\bf Q}_{1}^{-1}({\bf r}_{2}-{\bf g}_{1}\tilde{m}_{1})+(m_{1}-\tilde{m}_{1})^{2}{\bf g}_{1}^{\top}{\bf Q}_{1}^{-1}{\bf g}_{1},$ (33)
by defining
$\tilde{m}_{1}=\frac{{\bf g}_{1}^{\top}{\bf Q}_{1}^{-1}}{{\bf g}_{1}^{\top}{\bf Q}_{1}^{-1}{\bf g}_{1}}{\bf r}_{2}:={\bf w}_{1}^{\top}{\bf r}_{2}\in\mathbb{R}.$ (34)
Note that $\tilde{m}_{1}$ in (34) is the MVU estimate of $m_{1}$, and
${\bf w}_{1}=\frac{{\bf Q}_{1}^{-1}{\bf g}_{1}}{{\bf g}_{1}^{\top}{\bf Q}_{1}^{-1}{\bf g}_{1}}$ (35)
is the MVU estimator as a combining vector. Using (33), we rewrite (32) as
$\hat{m}_{1}=\underset{m_{1}\in\mathcal{M}_{1}}{\arg\min}\;\;(m_{1}-\tilde{m}_{1})^{2}.$ (36)
The optimal decision for User $2$ is to select $m_{1}\in\mathcal{M}_{1}$ that is closest to the MVU estimate $\tilde{m}_{1}\in\mathbb{R}$.
Next, we formulate the receive SNR in estimating $m_{1}$ at User 2. By putting ${\bf r}_{2}$ (in (29)) into (34), we rewrite (34) as
$\tilde{m}_{1}={\bf w}_{1}^{\top}{\bf g}_{1}m_{1}+{\bf w}_{1}^{\top}\big{(}({\bf I}+{\bf F}_{1}{\bf F}_{2}){\bf n}_{1}+{\bf F}_{1}{\bf n}_{2}\big{)}.$ (37)
Since $\tilde{m}_{1}$ is also a sufficient statistic for estimating $m_{1}$, by using (37), we can define the receive SNR in estimating $m_{1}$ as
@@
We note that the combining vectors for decoding in (25) and (35) are represented as functions of the encoding schemes of the users, i.e., ${\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, and ${\bf F}_{2}$. By plugging (25) in (28) and (35) in (38), we obtain the SNR expression that only depends on the encoding schemes, given by
@@
Thus, the joint design of the encoding and decoding schemes is equivalently reduced to the design of only the encoding schemes with the derived form of SNRs in (39).
@@ III-C Conversion of Optimization Problem
subject to $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}\leq NP,\quad\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\leq NP,$ (41)
where $\text{SNR}_{i}$ and ${\bf x}_{i}$, $i\in\{1,2\}$, are functions of the variables, ${\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, and ${\bf F}_{2}$, and the expectation is taken over the distributions of the random variables, ${m}_{1}$, ${m}_{2}$, ${\bf n}_{1}$, and ${\bf n}_{2}$. However, it is intractable to directly solve the optimization problem in (41) due to (i) the coupling between the encoding schemes and the SNR expression in (39) and (ii) the power constraints in (41). We thus propose to transform it into a more tractable optimization problem through several steps as explained below. We start by stating the following remark.
@@
The sum-error metric, $\mathcal{E}_{1}(\text{SNR}_{1})+\mathcal{E}_{2}(\text{SNR}_{2})$, is a decreasing function of $P$.
We further discuss the validity of Remark 2 within our linear coding framework in Appendix D. Based on the relationship between sum-error and power in Remark 2, we introduce an alternative optimization problem to (41), which aims to minimize the maximum transmit power, given by
$\underset{{\bf g}_{1},{\bf F}_{1},{\bf g}_{2},{\bf F}_{2},\bar{P}}{\text{minimize}}$ $\bar{P}$
@@
$\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}\leq N\bar{P},\quad\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\leq N\bar{P},$ (42)
where the auxiliary variable $\bar{P}$ is introduced. Note that the optimal solutions for (41) can be obtained by solving (42) through adjusting the value of $w$ based on Remark 2; if the obtained objective value $\bar{P}$ in (42) is smaller than $P$, we keep decreasing the value of $\omega$ and solve the above problem until $\bar{P}$ reaches $P$. However, solving (42) under the sum-error equality constraint involves identifying the optimal pair of $\text{SNR}_{1}$ and $\text{SNR}_{2}$, since multiple pairs may satisfy the constraint. To mitigate the optimization complexity, we consider individual constraints on $\text{SNR}_{1}$ and $\text{SNR}_{2}$ for optimization rather than the sum-error constraint. We subsequently reformulate (42) and obtain
$\underset{{\bf g}_{1},{\bf F}_{1},{\bf g}_{2},{\bf F}_{2},\bar{P}}{\text{minimize}}$ $\bar{P}$
subject to $\text{SNR}_{1}=\eta_{1},\quad\text{SNR}_{2}=\eta_{2},$
$\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}\leq N\bar{P},\quad\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\leq N\bar{P},$ (43)
where $\eta_{i}\in\mathbb{R}^{+}$ is the target SNR for the message $m_{i}$, $i\in\{1,2\}$. In (43), a feasible solution for $({\bf g}_{1},{\bf F}_{1},{\bf g}_{2},{\bf F}_{2})$ always exists for any $\eta_{1},\eta_{2}\in\mathbb{R}^{+}$ because the SNR constraints can be satisfied by adjusting the magnitudes of ${\bf g}_{1}$ and ${\bf g}_{2}$ in (39).
If $\bar{P}$ is an optimal value of the problem in (43), at least one of the power constraints in (43) should be satisfied in equality, i.e., $\bar{P}=\max\{\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]},\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\}/N$. If $\bar{P}\neq\max\{\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]},$ $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\}/N$, $\bar{P}$ is not an optimal solution since $\max\{\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}$ $\big{[}\|{\bf x}_{1}\|^{2}\big{]}$, $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\}/N$ can be the smaller objective value in (43). This insight leads us to reformulate (43) equivalently to
$\underset{{\bf g}_{1},{\bf F}_{1},{\bf g}_{2},{\bf F}_{2}}{\text{minimize}}$ ${\max}\big{\{}\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]},\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\big{\}}$
@@
We show, in the following lemma, that the optimal solutions of the original sum-error minimization problem with the power constraints in (41) can be obtained by solving the max-power minimization problem with the SNR constraints in (44).